# Bayesian Uncertainty Quantification and Reliability Assessment for Mars Sample Return

### v1

@@ 3.4 Cross validation in a Bayesian setting
Now that we have discussed all the elements that are used for generating the various posterior distributions, we move to the final step which is the simulation of the $\text{P}_{CNA}$.
@@
3.5 Simulation of probability of containment not assured $(\text{P}_{CNA})$
Our end goal is to calculate the probability that the peak OS acceleration exceeds 3000 Gs, we will call this $z_{crit}$. In order to do this, we run a simulation based on Algorithm 3, fixing large $N$ and $M$. For each iteration of the simulation, we sample from the posterior distributions of parameters for our Weibull and Normal variables respectively in order to then generate a new set of input values $\mathbf{s}_{0}$. Additionally, we sample from the posterior distributions of the spatial range parameters, $\theta_{1:K}$ as input for the UQ model. Finally, we calculate the exceedance probability based on our previous GP assumption that $Z\sim N_{n}(X\beta,\Sigma)$. What results is a sample of probabilities that can be treated as the posterior distribution of the probability of exceeding 3000 Gs.
@@ 3.5 Simulation of probability of containment not assured $(\text{P}_{CNA})$
Algorithm 3 $\text{P}_{CNA}$ Simulation
@@
4 Application and Results
The overall analysis was completed in several steps, the first of which was to estimate the parameters for the Weibull and Normal distributions of the respective input variables. Next we estimated the spatial range parameters for the spatial model as well as evaluated the performance of our UQ model. Finally, we used Algorithm 3 to compute the distribution of the $\text{P}_{CNA}$.
4.1 Input variable analysis
Given the distributional assumptions on our 15 variables, we find the MLE values for $\mu$, $\sigma$ and $\alpha$, $\beta$, respectively; additionally, we apply the AM algorithm using all three of the aforementioned prior distributions (dependent on the distributional assumption). The MLE values give us an initial estimate on the parameters of the distributions. Allowing for different priors indicates how influential the choice of prior is on the posterior distribution. Additionally, when comparing the posterior means to the MLE values we can check the robustness of the results.
The results for the normally distributed variables can be seen in Figure 1 where the mean of the posterior distributions along with their 95% credible intervals (CIs) are plotted; additionally, the MLE s and their 95% confidence intervals are included for comparison. As observed in Figure 1(a), all the means and CIs are very similar. In Figure 1(b) we see less agreement in the scale parameter values, with some variation in the mean values as well as differing lengths of CIs; however, overall, the agreement across the priors is fairly consistent with no statistically significant differences.
@@ 4.1 Input variable analysis
4.2 Spatial range parameters analysis
Applying the GP approach to our data, we have $\mathbf{X}=(X0001,\ldots,X0015)$ as our input variables with the previously generated LHS as our input values for those variables. For our output variable $Z$, or the peak OS acceleration, we use the 25 realizations from JPL’s work. In our analysis, for numerical stability, we re-scaled the output data by dividing all output values by a factor of 1000, making our problem determining the probability that the peak OS acceleration is greater than $3.0$ when the input variables are random.
For simplicity, we assume a constant mean response for $X\beta$. Initially, we seek to determine the MLE values for the spatial range parameters $\theta_{1:K}$ with the overarching goal of generating posterior distributions for $\theta_{1:K}$ and then drawing from those distributions during the prediction (kriging) process. Considering several values of $\lambda>0$ via CV, we choose the value of $\lambda=2$ for our penalty term. We also completed the kriging step within the CV analysis at this point to see how our model performs in predicting the peak OS acceleration. The CV plot as well as a diagnostic plot of the model’s predictive quality can be seen in Figure 3.
From Figure 3, we see that $\lambda=2$ does provide the smallest CV value. In Figure 3 we see that the variance of the noise is greater than that of the signal. From this we consider that it is possible that the range of input values is too narrow; that is, we would get greater signal to noise ratio with a wider range of inputs. We also note that there are no simulated values in the critical region; that is, our $Z_{1:25}$ range from $2.474$ to $2.749$, so we have no data over $3.0$. Furthermore, kriging with a constant mean is a form of interpolation, which means it will never produce a predicted value outside the range of data and input values far from the test data set will simply be predicted back to the overall mean.
In spite of this critique, which we will discuss further in the next section, we move forward with the Bayesian analysis. We use Equation 3 with $\theta_{1:K}\sim N(\tau,\nu^{2})$ as the prior for the spatial range parameters; however, we must first determine appropriate values for $\tau$ and $\nu^{2}$. We start by using an Empirical Bayes approach, considering the MLE values of the Normal prior distribution for the $\theta_{i}$’s which are $\hat{\tau}=\bar{\theta}$ and $\hat{\nu}^{2}=\frac{1}{K}\sum_{i=1}^{K}{(\theta_{i}-\hat{\tau})^{2}}$. Here, as we do not have the actual values of $\theta_{1:K}$, we use the values generated from Equation 2 with $\lambda=2$ to calculate $\hat{\tau}$ and $\hat{\nu}^{2}$. However, when calculating $\hat{\nu}^{2}$ we get a value that is too small; that is, the MLE value of $\hat{\nu}^{2}$ suggests less variability than we believe truly exists. Therefore, we turn to the output of the Hessian matrix, $H$, from the optimization of Equation 2 and use $\hat{\nu}^{2}=\frac{1}{K}\sum_{i=1}^{K}{H_{i,i}^{-1}}$ as our estimate for $\nu^{2}$. This value provides a much more practical estimate of the variance at $0.259$ as compared to the MLE of $0.0364$.
@@ 4.2 Spatial range parameters analysis
(d) Observed vs. Expected peak OS acceleration where observed are calculated using kriging methods described in Section 3.3
@@
After making the adjustment for $\hat{\nu}^{2}$, we then turn to CV to determine if the value of $\hat{\tau}=1.49$ as calculated from the output of optimizing Equation 2 is the best choice. In order to do this, for each potential value of $\tau$ under consideration, we run the adaptive algorithm $n=25$ times (as we used leave one out CV). Because of this, we narrowed the CV options to five potential values, $\tau=\{1.00,1.10,1.25,1.49,1.75\}$ based on some initial runs of the CV with fewer iterations for the AM algorithm. We see the results of the more extensive (in terms of number of iteration in the AM algorithm) CV analysis in Figure 4a which shows that $\hat{\tau}=1.25$ is the best option for our estimate of the mean of our prior distribution. To check the appropriateness of the choice for $\nu^{2}$, we did an additional round of CV with the same values of $\tau$ while using $\nu^{2}=0.5$. The CV scores with $\nu^{2}=0.5$ were consistently larger than when using $\nu^{2}=0.259$.
@@
Once we determined the best choices for the hyperparameters of the prior distribution $(\tau=1.25,\nu^{2}=0.259)$, we examine the model performance using the chosen parameters. We also compare these results to those using the regularized REML approach with $\lambda=2$ as the penalty term. Comparing Figures 3 and 4, we see that the prediction using the Bayesian approach is slightly worse; however, this is to be expected given that we have allowed for more uncertainty within the choice of the spatial range parameters.
@@
Finally, we examine the resulting values of the spatial range parameters from the Bayesian analysis. In Figure 5, we see the MLE results from Equation 2 plotted with their 95% confidence intervals as well as the posterior means and 95% CIs produced by the AM algorithm. As we can see, the two methods are fairly comparable, with the posterior mean estimates from the Bayesian method pooling toward the common mean, driven by the choice of $\tau=1.25$ as the mean of the prior distribution. This regularization in the Bayesian context is to be expected given the few number of data points we have, making the influence of the prior stronger.
@@
(b) Observed vs. Expected peak OS acceleration where observed are calculated using kriging methods described in Section 3.3
@@
Figure 5: Comparison of MLE and Bayesian results
@@
4.3 Final simulation results
@@
We run the end-to-end simulation found in Algorithm 3 using two settings. In both settings A and B, we use the posterior distributions for $\alpha,\beta,\mu,\sigma^{2}$ generated by applying the AM algorithm with the respective Jeffreys’ priors. For the spatial range parameters $\theta_{1:15}$ in Setting A, we use the MLE results given by Equation 2 with penalty term $\lambda=2$. And in Setting B, for the spatial range parameters $\theta_{1:15}$ we use the posteriors produced by applying the AM algorithm with Equation 3 and prior $\theta_{k}\sim N(1.25,0.259)$. In both settings, we let $N=2000$ and $M=1000$. The resulting posterior distributions of the $\text{P}_{CNA}$ can be seen in Figure 6.
@@
As observed in Figure 6a under Setting A, both the median and mean of the posterior distribution fail to meet the target value of one, or the probability of one in a million. However, in Figure 6b, the median of the posterior is well below the target value at $0.13$ and the mean only slightly exceeds the target value at $1.16$. We also see a narrower credible interval under Setting B. Thus, even though we have included a greater amount of uncertainty by varying the values of the spatial range parameters in Setting B, the $\text{P}_{CNA}$ appears to be reduced. A deeper dive into this result concluded that this change is due to the slight decrease in the spatial range parameter estimates and is specifically tied to the choice of prior parameters.
@@
While this raises some concerns, we emphasize the extensive exploration in selecting the parameter values of the prior via CV. That is, in the context of this problem, we have incorporated empirical evidence by beginning with the MLE values for the spatial range parameters and then applying CV to choose hyperparameter values for the prior on the spatial range parameters in the Bayesian framework. Although the variability in $\theta_{1:K}$ has a strong influence on the final $\text{P}_{CNA}$ distribution, we have built some assurance in the results in Figure 6b with the systematic approaches outlined above. The implications of these findings and potential avenues for future research are discussed in the following section.
(a) Plot of posterior distribution of the $\text{P}_{CNA}$ under Setting A
(b) Plot of posterior distribution of the $\text{P}_{CNA}$ under Setting B
Figure 6: Comparison of Posterior Distributions of $\text{P}_{CNA}$ between settings A and B
5 Conclusion and Future Work
@@
The general process outlined and applied above gives a Bayesian approach to UQ of computer simulations. In context of the overarching goal for the MSR program, we have demonstrated a model for providing a measurement of the $\text{P}_{CNA}$ for the EES. By employing Bayesian methods, we produce a probability density of the $\text{P}_{CNA}$ as opposed to a singular point estimate. While our method and the subsequent results have limitations, we have ascertained that the predominant goal of ensuring containment at least equal to $99.9999\%$ is conceivable for this particular component. Additionally, providing decision makers with CIs allows for further risk assessment and analysis in the event that the stated goal has some flexibility. Indeed, these results would be factored into larger reliability analyses (Cataldo et al., nd, ) as part of the assurance case MSR is developing per NASA’s requirements (NASA Office of Safety and Mission Assurance,, 2021).
@@
Several of the limitations within this analysis suggest directions for future projects. The first was mentioned in Section 4.2 in regard to kriging with a constant mean. While this element of the model could be modified, we believe the real issue here is the data used to generate the initial set of output values. We mentioned the engineers running the experiment used LHS when sampling from the input variables. This is a perfectly valid method of sampling as outlined by McKay et al., (1979). However, the 25 LHS samples fail to produce an example of the peak OS acceleration exceeding 3000 Gs when run through the chosen computer model. Therefore, we ask, if the simulation does not produce a critical value, how can we understand the conditions under which a critical value might occur in reality? To avoid this issue, we suggest a study on the comparison of sampling methods in respect to their impact on the $\text{P}_{CNA}$.
@@
Although we believe that our method effectively exploits the data that are available, we also believe it highlights the disadvantages of working with such a limited dataset. As the initial computer model only produced 25 simulations of the peak OS acceleration, we were very limited in testing the robustness of our UQ model. There were multiple reasons for the number of runs conducted by JPL which include the cost of experiments as well as the length of time the simulation software takes. However, given the work we have done here, our counterparts at JPL have been receptive to the following suggestion. We propose the engineers generate more simulations using their computer model, perhaps using a sampling scheme as indicated by our next planned study. This would allow us to reapply the methods used in this paper and explore further alternative techniques if we meet the same sensitivity issues in regards to the influence of the choice of spatial range parameters in the GP model.
@@
Finally, we outlined the chain of influence the choice of the spatial range parameters has on the final probability distribution. To explore this influence further, the future study with the incorporation of additional data should consider alterations to the model. These alterations would include examining a change to the GP’s covariance structure, and/or a change to the GP’s mean function, as well as a larger scale change in the distributional assumptions for the output variable $Z$ (e.g., a t-distribution). Exploring these different avenues, bolstered by additional data, will allow us to hone our methodology within this particular problem set and strengthen our abilities to apply such methods to additional aspects of the MSR program.
@@
In conclusion, our work shines a light on the significant potential of the Bayesian approach in UQ for computer simulations, particularly in the context of complex projects such as the MSR. The highlighted limitations and subsequent recommendations not only provide a road map for immediate improvements but also emphasize the interdisciplinary nature of this research, bridging the gap between statistical methodologies and engineering challenges. As we advance further into an era dominated by simulations and computational models, understanding the underpinnings of their uncertainties becomes paramount. As we move forward, augmenting our dataset and refining our methodologies will be instrumental in enhancing the reliability and precision of our computational assessments.
Supplementary Materials
In the supplementary materials we define the settings we used when applying the AM algorithm. We also include information on hyperparameter choices and further algorithm settings when using CV for the spatial range parameters. In separate files, we include the data and code.
@@
SUPPLEMENTARY MATERIAL
Data:
The description of the different variables as well as the number of observations and the distributional assumptions can be seen in Table 2.
X0001 $IM7-E_{a},E_{b},E_{c}$ Young’s Modulus 8 Normal
X0002 $IM7-G_{ab},G_{bc},G_{ca}$ Shear Modulus 4 Normal
X0003 $IM7-X_{c}$ Compressive Strength (0 Deg) 3 Weibull
X0004 $IM7-X_{t}$ Tensile Strength (0 Deg) 4 Weibull
X0005 $IM7-Y_{c}$ Compressive Strength (90 Deg) 3 Weibull
X0006 $IM7-Y_{t}$ Tensile Strength (90 Deg) 4 Weibull
X0007 $IM7-S_{c}$ Shear Strength 4 Weibull
X0008 $Kv-E_{a},E_{c}$ Young’s Modulus (0 Deg) 4 Normal
X0009 $Kv-E_{b}$ Young’s Modulus (90 Deg) 4 Normal
X0010 $Kv-G_{ab},G_{bc},G_{ca}$ Shear Modulus 4 Normal
X0011 $Kv-X_{c}$ Compressive Strength (0 Deg) 12 Weibull
X0012 $Kv-X_{t}$ Tensile Strength (0 Deg) 4 Weibull
X0013 $Kv-Y_{c}$ Compressive Strength (90 Deg) 11 Weibull
X0014 $Kv-Y_{t}$ Tensile Strength (90 Deg) 4 Weibull
X0015 $Kv-S_{c}$ Shear Strength 4 Weibull
Table 2: Description of input variables
AM algorithm settings:
We refer the reader to Haario et al., (2001) for an in-depth description of the AM algorithm. For every application of the AM algorithm unless otherwise noted, we set the total number of simulations to: $t=100,000$; we also set the following values: $s_{d}=\frac{(2.4)^{2}}{d}$ ($d$ is the dimension of the sample parameter space), $\epsilon=0.0001$, $t_{0}=10,000$ ($t_{0}$ is the length of the initial period (Haario et al.,, 2001), $t_{1}=10$ ($t_{1}$ is length of iterations between updating the covariance matrix), as well as $t_{2}=100$ ($t_{2}$ is the interval between updates of our output parameter matrix). Therefore, we end up with an output parameter matrix of dimension $\frac{t}{t_{2}}\times d=1,000\times d$. The number of total simulations was chosen after some initial runs at varying values; the choice of $t=100,000$ was based on the trace plots and running average plots. The values for $t_{0},t_{1}$, and $t_{2}$ were based on the speed of the algorithm as well as our confidence in the initial covariance matrix (in regards to $t_{0}$). After establishing the above settings of the AM algorithm, we looked at trace plots and running average plots as well as the GCD for each application of the AM algorithm in order to assess the burn-in rate for the algorithm. Taking into consideration all of these diagnostics, we removed the first 20% of the runs as burn-in for each use of the AM algorithm unless otherwise specified.
Hyperparameter choices for input variables:
For the Normally distributed variables, the hyperparameters for the Normal Inverse Gamma (NIG) were tuned and set to $m=10$, $c=0.003$, $a=0.1$, and $b=0.5$. And for the Weibull variables, the hyperparameters for the Inverse Gamma (IG) prior were tuned and set the hyperparameters to $\gamma=1,\zeta=2$.
Spatial range parameter CV:
We performed the CV using the following two settings for the AM algorithm, Setting 1: $t=5,000,t_{0}=500,t_{1}=10,t_{2}=50$ and Setting 2: $t=10,000,t_{0}=1,000,t_{1}=10,t_{2}=100$. We did not increase the total number of iterations to $100,000$ due to the projected length of time the CV would take. In both CV runs with the above settings, the conclusions agreed with the results presented in the body of the paper.

### v2

@@ 3.1.4   Cross validation in a Bayesian setting
Now that we have discussed all the elements that are used for generating the various posterior distributions, we move to the final step which is the simulation of the $\text{P}_{f}$.
@@
3.1.5   Simulation of probability of failure $(\text{P}_{f})$
Our end goal is to calculate the probability that the peak acceleration exceeds 3000 Gs, we will call this $z_{crit}$. In order to do this, we run a simulation based on Algorithm 3, fixing large $N$ and $M$. For each iteration of the simulation, we sample from the posterior distributions of parameters for our Weibull and Normal variables respectively in order to then generate a new set of input values $\mathbf{s}_{0}$. Additionally, we sample from the posterior distributions of the spatial range parameters, $\theta_{1:K}$ as input for the UQ model. Finally, we calculate the exceedance probability based on our previous GP assumption that $Z\sim N_{n}(X\beta,\Sigma)$. What results is a sample of probabilities that can be treated as the posterior distribution of the probability of exceeding 3000 Gs.
@@ 3.1.5   Simulation of probability of failure $(\text{P}_{f})$
Algorithm 3 $\text{P}_{f}$ Simulation
@@
3.2   Application and Results
The overall analysis was completed in several steps, the first of which was to estimate the parameters for the Weibull and Normal distributions of the respective input variables. Next we estimated the spatial range parameters for the spatial model as well as evaluated the performance of our UQ model. Finally, we used Algorithm 3 to compute the distribution of the $\text{P}_{f}$. All analysis was completed using R Software (R Core Team,, 2022).
3.2.1   Input variable analysis
Given the distributional assumptions on our 15 variables, we find the REML values for $\mu$, $\sigma$ and $\alpha$, $\beta$, respectively; additionally, we apply the AM algorithm using all three of the aforementioned prior distributions (dependent on the distributional assumption). The REML values give us an initial estimate on the parameters of the distributions. Allowing for different priors indicates how influential the choice of prior is on the posterior distribution. Additionally, when comparing the posterior means to the REML values we can check the robustness of the results.
The following description of the settings used for the AM algorithm follow the notation provide by Haario et al., (2001). For every application of the AM algorithm unless otherwise noted, we set the total number of simulations to: $t=100,000$; we also set the following values: $s_{d}=\frac{(2.4)^{2}}{d}$ (where $d$ is the dimension of the sample parameter space), $\epsilon=0.0001$, $t_{0}=10,000$, $t_{1}=10$, as well as $t_{2}=100$ where $t_{2}$ is the interval between updates of our output parameter matrix. Therefore, we end up with an output parameter matrix of dimension $\frac{t}{t_{2}}\times d=1,000\times d$. The number of total simulations was chosen after some initial runs at varying values; the choice of $t=100,000$ was based on the trace plots and running average plots. The values for $t_{0},t_{1}$, and $t_{2}$ were based on the speed of the algorithm as well as our confidence in the initial covariance matrix (in regards to $t_{0}$). After establishing the above settings of the AM algorithm, we looked at trace plots and running average plots as well as the GCD for each application of the AM algorithm in order to assess the burn-in rate for the algorithm. Taking into consideration all of these diagnostics, we removed the first 20% of the runs as burn-in for each use of the AM algorithm unless otherwise specified.
The results for the normally distributed variables can be seen in Figure 1 where the mean of the posterior distributions along with their 95% credible intervals (CIs) are plotted; additionally, the REML s and their 95% confidence intervals are included for comparison. As observed in Figure 1(a), all the means and CIs are very similar. In Figure 1(b) we see less agreement in the scale parameter values, with some variation in the mean values as well as differing lengths of CIs; however, overall, the agreement across the priors is fairly consistent with no statistically significant differences.
@@ 3.2.1   Input variable analysis
3.2.2   Spatial range parameters analysis
Applying the GP approach to our data, we have $\mathbf{X}=(X0001,\ldots,X0015)$ as our input variables with the previously generated LHS as our input values for those variables. For our output variable $Z$, or the peak acceleration, we use the 25 realizations from Jet Propulsion Laboratory (JPL)’s work. In our analysis, for numerical stability, we re-scaled the output data by dividing all output values by a factor of 1000, making our problem determining the probability that the peak acceleration is greater than $3.0$ when the input variables are random.
For simplicity, we assume a constant mean response for $X\beta$ (e.g. $\mathbf{X}=\mathbf{1}_{n\times 1}$). Initially, we seek to determine the REML values for the spatial range parameters $\theta_{1:K}$ with the overarching goal of generating posterior distributions for $\theta_{1:K}$ and then drawing from those distributions during the prediction (kriging) process. Considering several values of $\lambda>0$ via CV, we choose the value of $\lambda=2$ for our penalty term. We also completed the kriging step within the CV analysis at this point to see how our model performs in predicting the peak acceleration. The CV plot as well as a diagnostic plot of the model’s predictive quality can be seen in Figure 3.
From Figure 3, we see that $\lambda=2$ does provide the smallest CV value. In Figure 3 we see that the variance of the noise is greater than that of the signal. This is supported by the weak correlation of $0.393$ and the low signal to noise ratio of $0.390$. From this we consider that it is possible that the range of input values is too narrow; that is, we would get greater signal to noise ratio with a wider range of inputs. We also note that there are no simulated values in the critical region; that is, our $Z_{1:25}$ range from $2.474$ to $2.749$, so we have no data over $3.0$. Furthermore, kriging with a constant mean is a form of interpolation, which means it will never produce a predicted value outside the range of data and input values far from the test data set will simply be predicted back to the overall mean.
@@ 3.2.2   Spatial range parameters analysis
(d) Observed vs. Expected peak acceleration where observed are calculated using kriging methods described in Section 3.1.3
@@
In spite of this critique, which we discuss further in the next section, we move forward with the Bayesian analysis. We use Equation 6 with $\theta_{1:K}\sim N(\tau,\nu^{2})$ as the prior for the spatial range parameters; however, we must first determine appropriate values for $\tau$ and $\nu^{2}$. We start by using an Empirical Bayes approach, considering the REML values of the Normal prior distribution for the $\theta_{i}$’s which are $\hat{\tau}=\bar{\theta}$ and $\hat{\nu}^{2}=\frac{1}{K}\sum_{i=1}^{K}{(\theta_{i}-\hat{\tau})^{2}}$. Here, as we do not have the actual values of $\theta_{1:K}$, we use the values generated from Equation 5 with $\lambda=2$ to calculate $\hat{\tau}$ and $\hat{\nu}^{2}$. This approach suffices for the estimation of $\hat{\tau}=1.49$, as the REML values of the $\theta_{i}$’s provide unbiased point estimates of their central tendency; however, the REML-derived value of $\hat{\nu}^{2}$ underestimates the true variability in the $\theta_{i}$’s due to its reliance solely on the spread of point estimates without accounting for uncertainty in parameter estimation. This underestimation may be particularly severe here given the small sample size. To address this, we instead use the output of the Hessian matrix, $H$, from the optimization of Equation 5 and use $\hat{\nu}^{2}=\frac{1}{K}\sum_{i=1}^{K}{H_{i,i}^{-1}}$ as our estimate for $\nu^{2}$. This estimation accounts for the curvature of the likelihood function and better represents the uncertainty in the spatial range parameters. Using this method, we obtain $\hat{\nu}^{2}=0.259$, a more realistic value that incorporates the variability in both the estimated parameters and the likelihood surface itself.
@@
After making the adjustment for $\hat{\nu}^{2}$, we then turn to CV to determine if the MLE derived value of $\hat{\tau}=1.49$ as calculated from the output of optimizing Equation 5 is the best choice. In order to do this, for each potential value of $\tau$ under consideration, we run the adaptive algorithm $n=25$ times (as we used leave one out CV). Because of the computing time necessary to implement leave one out CV, we narrowed the CV options to five potential values, $\tau=\{1.00,1.10,1.25,1.49,1.75\}$ based on some initial runs of the CV with fewer iterations for the AM algorithm. We see the results of the more extensive (in terms of number of iteration in the AM algorithm) CV analysis in Figure 4a which shows that $\hat{\tau}=1.25$ is the best option for our estimate of the mean of our prior distribution. To check the appropriateness of the choice for $\nu^{2}$, we did an additional round of CV with the same values of $\tau$ while using $\nu^{2}=0.5$. The CV scores with $\nu^{2}=0.5$ were consistently larger than when using $\nu^{2}=0.259$. Because of the amount of computing time necessary to run the CV, we did not examine different combinations of $\tau$ and $\nu^{2}$ simultaneously.
@@
Once we determined the best choices for the hyperparameters of the prior distribution $(\tau=1.25,\nu^{2}=0.259)$, we examine the model performance using the chosen parameters. We also compare these results to those using the regularized REML approach with $\lambda=2$ as the penalty term. Comparing the reported correlation values in Figures 3 and 4 ($0.393$ and $0.318$, respectively), we see that the prediction using the Bayesian approach is slightly worse; however, this is to be expected given that we have allowed for more uncertainty within the choice of the spatial range parameters.
@@
Finally, we examine the resulting values of the spatial range parameters from the Bayesian analysis. In Figure 5, we see the REML results from Equation 5 plotted with their 95% confidence intervals as well as the posterior means and 95% CIs produced by the AM algorithm. As we can see, the two methods are fairly comparable, with the posterior mean estimates from the Bayesian method pooling toward the common mean, driven by the choice of $\tau=1.25$ as the mean of the prior distribution. This regularization in the Bayesian context is to be expected given the few number of data points we have, making the influence of the prior stronger.
@@
(b) Observed vs. Expected peak acceleration where observed are calculated using kriging methods described in Section 3.1.3
@@
Figure 5: Comparison of REML and Bayesian results
@@
3.2.3   Final simulation results
@@
We run the end-to-end simulation found in Algorithm 3 using two settings. In both settings A and B, we use the posterior distributions for $\alpha,\beta,\mu,\sigma^{2}$ generated by applying the AM algorithm with the respective Jeffreys’ priors. For the spatial range parameters $\theta_{1:15}$ in Setting A, we use the REML results given by Equation 5 with penalty term $\lambda=2$. And in Setting B, for the spatial range parameters $\theta_{1:15}$ we use the posteriors produced by applying the AM algorithm with Equation 6 and prior $\theta_{k}\sim N(1.25,0.259)$. In both settings, we let $N=2000$ and $M=1000$. The resulting posterior distributions of the $\text{P}_{f}$ can be seen in Figure 6.
@@
As observed in Figure 6a under Setting A, both the median and mean of the posterior distribution fail to meet the target value of one, or the probability of one in a million. However, in Figure 6b, the median of the posterior is well below the target value at $0.13$ and the mean only slightly exceeds the target value at $1.16$. We also see a narrower credible interval under Setting B. Thus, even though we have included a greater amount of uncertainty by varying the values of the spatial range parameters in Setting B, the $\text{P}_{f}$ appears to be reduced. A deeper dive into this result concluded that this change is due to the slight decrease in the spatial range parameter estimates and is specifically tied to the choice of prior parameters.
@@
While this raises some concerns, we emphasize the extensive exploration in selecting the parameter values of the prior via CV. That is, in the context of this problem, we have incorporated empirical evidence by beginning with the REML values for the spatial range parameters and then applying CV to choose hyperparameter values for the prior on the spatial range parameters in the Bayesian framework. Although the variability in $\theta_{1:K}$ has a strong influence on the final $\text{P}_{f}$ distribution, we have built some assurance in the results in Figure 6b with the systematic approaches outlined above. The implications of these findings and potential avenues for future research are discussed in the following section.
(a) Plot of posterior distribution of the $\text{P}_{f}$ under Setting A
(b) Plot of posterior distribution of the $\text{P}_{f}$ under Setting B
Figure 6: Comparison of Posterior Distributions of $\text{P}_{f}$ between settings A and B
4 Advice to practitioners
@@
The general process outlined and applied above gives a Bayesian approach to UQ of computer simulations. Using the simulation of peak acceleration for a space sample return mission as a case study, we have demonstrated a model for providing a measurement of the $\text{P}_{f}$ for an Earth reentry capsule. By employing Bayesian methods, we have produced a probability density of the $\text{P}_{f}$ as opposed to a singular point estimate. While our method and the subsequent results have the limitations indicated below, our analysis suggests that the predominant goal of ensuring a safe landing at least equal to $99.9999\%$ may be conceivable for this particular component of a reentry capsule. Additionally, providing decision makers with CIs allows for further risk assessment and analysis in the event that the stated goal has some flexibility (Cataldo et al., 2025b, ).
@@
Several of the limitations within this analysis suggest directions for future projects. The first was mentioned in Section 3.2.2 in regard to kriging with a constant mean. While this element of the model could be modified, we believe the real issue here is the data used to generate the initial set of output values. We mentioned the engineers running the experiment used LHS when sampling from the input variables. This is a perfectly valid method of sampling as outlined by McKay et al., (1979). However, the 25 LHS samples fail to produce an example of the peak acceleration exceeding 3000 Gs when run through the chosen computer model. Although the experiment was constructed in line with well-established principles for LHS sampling, it is nevertheless problematic that we are trying to calculate an exceedance probability for 3000 Gs when no member of the simulation exceeds that level. To avoid this issue, we suggest a study on the comparison of sampling methods in respect to their impact on the $\text{P}_{f}$.
@@
Although we believe that our method effectively exploits the data that are available, we also believe it highlights the disadvantages of working with such a limited dataset. As the initial computer model only produced 25 simulations of the peak acceleration, we were very limited in testing the robustness of our UQ model. There were multiple reasons for the number of runs conducted by JPL, which include the cost of experiments as well as the length of time the simulation software takes. However, given the work we have done here, our counterparts at JPL have been receptive to the following suggestion. We propose the engineers generate more simulations using their computer model, perhaps using a sampling scheme as indicated by our next planned study. This would allow us to reapply the methods used in this paper and explore further alternative techniques if we meet the same sensitivity issues in regards to the influence of the choice of spatial range parameters in the GP model.
@@
Finally, we outlined the chain of influence the choice of the spatial range parameters has on the final probability distribution. To explore this influence further, the future study with the incorporation of additional data should consider alterations to the model. These alterations would include examining a change to the GP’s covariance structure, and/or a change to the GP’s mean function, as well as a larger scale change in the distributional assumptions for the output variable $Z$ (e.g., a t-distribution). Exploring these different avenues, bolstered by additional data, will allow us to hone our methodology within this particular problem set and strengthen our abilities to apply such methods to similar problems in other space missions.
@@
In conclusion, our work shines a light on the significant potential of the Bayesian approach in UQ for computer simulations, particularly in the context of complex projects such as, for example, Mars Sample Return (MSR) (Cataldo et al.,, 2024; Sarli et al.,, 2024). The highlighted limitations and subsequent recommendations not only provide a roadmap for immediate improvements but also emphasize the interdisciplinary nature of this research, bridging the gap between statistical methodologies and engineering challenges. As we advance further into an era dominated by simulations and computational models, understanding the underpinnings of their uncertainties becomes paramount. As we move forward, augmenting our dataset and refining our methodologies will be instrumental in enhancing the reliability and precision of our computational assessments.
@@
Acknowledgments
The authors would like to thank their colleagues Aaron Siddens and Kevin Carpenter at the Jet Propulsion Laboratory for providing the dataset used in this manuscript.
Funding
Funding provided by the National Aeronautics and Space Administration is gratefully acknowledged. Portions of the research were carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004).
Data Availability
Due to the proprietary nature of the original engineering simulations, the dataset used in this study cannot be publicly shared. However, a representative dataset that preserves the structure and statistical properties of the original data has been created to support reproducibility. This simulated dataset, along with annotated code used for the analysis, is available at https://doi.org/10.5281/zenodo.15855903.